### v1

# QGait: Toward Accurate Quantization for Gait Recognition with Binarized Input

Abstract
Existing deep learning methods have made significant progress in gait recognition. Typically, appearance-based models binarize inputs into silhouette sequences. However, mainstream quantization methods prioritize minimizing task loss over quantization error, which is detrimental to gait recognition with binarized inputs. Minor variations in silhouette sequences can be diminished in the network’s intermediate layers due to the accumulation of quantization errors. To address this, we propose a differentiable soft quantizer, which better simulates the gradient of the round function during backpropagation. This enables the network to learn from subtle input perturbations. However, our theoretical analysis and empirical studies reveal that directly applying the soft quantizer can hinder network convergence. We further refine the training strategy to ensure convergence while simulating quantization errors. Additionally, we visualize the distribution of outputs from different samples in the feature space and observe significant changes compared to the full precision network, which harms performance. Based on this, we propose an Inter-class Distance-guided Distillation (IDD) strategy to preserve the relative distance between the embeddings of samples with different labels. Extensive experiments validate the effectiveness of our approach, demonstrating state-of-the-art accuracy across various settings and datasets. The code will be made publicly available.
1 Introduction
In recent years, gait recognition has emerged as a prominent field due to its potential in long-range pedestrian retrieval based on walking patterns. Unlike other biometric technologies, gait, as a form of locomotion, is difficult to disguise [18; 8]. Additionally, gait recognition is robust against common covariates such as attire, carrying items, and standing conditions [13; 7]. Therefore, gait recognition methods hold significant promise for capturing biometric features remotely in uncontrolled environments, thus offering broad application prospects.
With the development of deep learning, gait recognition methods based on deep features [38; 27; 12; 15; 22] have achieved notable results on various datasets [40; 33; 43; 31; 46; 11; 10]. Currently, state-of-the-art gait recognition technologies can achieve recognition distances exceeding hundreds of meters even with 4K high-definition cameras [31]. However, existing methods have scarcely considered the increasing computational resources required by the advanced devices. In some scenarios, it may be necessary to test tens of thousands of subjects within a short timeframe. While methods based on CNNs [11; 43] or Transformers [10] have demonstrated excellent performance, edge devices in deployment environments cannot meet the demands of these resource-intensive algorithms. This necessitates reducing the memory and computational burden of gait recognition methods while preserving performance, enabling deployment on resource-constrained devices. Current efforts to compress and accelerate neural networks primarily focus on quantization [45; 6; 9; 19] and pruning [24; 44; 14]. Compared to pruning, quantization is more deployment-friendly and offers significant acceleration, particularly Quantization Aware Training [16].
Existing quantization methods [42; 1; 9] have achieved decent results on some tasks, but applying them to gait recognition tasks has failed to deliver comparable performance. These methods use a Straight-Through Estimator (STE) [2] to approximate the gradient of non-differentiable operations during the backward pass. While STE can smoothly transfer gradients, it is not sensitive to decimal-level changes. This may have no impact on other tasks but can be fatal for changes in gait posture. For example, in RGB input, the variation from 0 to 1 in a small area is not significant. However, for binarized gait silhouettes, this may indicate a change in posture. In this paper, we propose using parameterized soft quantization operators to prevent the network from accumulating quantization errors. Initially, we attempted to directly utilize a soft quantizer in the training process as a replacement for STE. However, through theoretical analysis and empirical validation, we have demonstrated that although this approach can simulate the quantization error introduced by non-differentiable operations, it significantly impacts the network’s convergence. To harness the strengths of both STE and the soft quantizer while mitigating their respective limitations, we propose a two-stage training strategy. In the first stage, we employ STE to facilitate the network’s convergence towards the vicinity of the optimal solution. Subsequently, we utilize the soft quantizer in the fine-tuning stage, enabling the simulation of decimal-level errors, which improves the network performance.
After quantization, the 8-bit models exhibit outstanding performance, surpassing even the full-precision model. However, when quantized to lower bit widths, the model’s accuracy experiences significant degradation. Through analysis of the sample distribution in feature space, we observed significant changes in the relative positions of samples between the 4-bit quantized model and the full-precision model. This change could be a significant contributing factor to the observed decrease in accuracy. This insight prompts us to consider whether knowledge distillation could be employed to transfer the knowledge from the full-precision model to the low-bit model. However, through an examination of the properties of traditional distillation [9; 17; 39], we found that such methods result in the low-bit model learning the numerical discrepancies between the outputs of the full-precision model and itself, which are discrepancies inherently introduced by quantization. Experimental results similarly demonstrate that this distillation approach does not yield benefits. Therefore, we propose an Inter-class Distance-guided Distillation method, allowing the network to focus more on the differences in sample distribution between classes rather than the numerical differences within classes. This distillation approach leads to a significant improvement in accuracy.
Our contributions can be summarized as follows:
We introduce, for the first time, a quantization-based compressed gait recognition network, which performs lossless compression and acceleration on existing methods as Fig. 1 shows.
We propose a two-stage quantization training strategy together with the IDD method, which simultaneously optimizes quantization error and task loss.
We conduct detailed experiments and the results demonstrate that our proposed QGait achieves comparable performance to full-precision networks with the lowest computational overhead.
2 Related Work
2.1 Gait Recognition
Gait recognition methods can generally be classified into two categories: model-based methods [20; 21] and appearance-based methods [37; 32]. Model-based approaches tend to utilize estimated human body models as input, such as 2D/3D poses and SMPL [26]. For instance, GaitGraph [34] employs graph convolutional networks for learning gait representations based on 2D skeletons, SMPLGait [43] leverages the 3D geometric information of SMPL models to enhance the learning of gait appearance features, and GaitTR [41] combines transformer and convolutional layers to represent spatial and temporal information respectively. However, although model-based methods theoretically exhibit robustness to factors like carrying items and clothing, they often perform poorly due to the error accumulation in the pre-processing stage while entailing high computational costs, thereby potentially lacking practicality in certain real-world scenarios. On the other hand, appearance-based methods directly learn shape features from input videos, offering simplicity and ease of use while preserving privacy. With the rise of deep learning, most current appearance-based works focus on spatial feature extraction and gait temporal modeling. Specifically, GaitSet [4] treats gait sequences as a set for the first time, employing the maximum function to compress sequences of frame-level spatial features. GaitPart [12] meticulously explores minor differences in various parts of the silhouette to model periodic gait features. GaitGL [22] addresses the limitations of spatially global gait representations in neglecting details, and locally based descriptors in failing to capture relationships between adjacent parts, by developing global and local convolutional layers. Building upon these works, GaitBase [11] significantly simplifies gait modeling and achieves excellent results with its simple yet effective design. Considering that existing methods often lack a practical perspective, it becomes necessary to design a lightweight network to aid gait recognition models in real-world applications. To the best of our knowledge, the proposed QGait is the first quantization-based compressed gait recognition network.
2.2 Network Quantization
Quantization [5; 3; 28] is a widely adopted technique in computer vision for compressing and accelerating neural networks. Depending on the mapping strategy, quantization can be categorized into uniform and non-uniform [25] quantization. Non-uniform quantization necessitates specific hardware support, thus current research predominantly focuses on uniform quantization exploration. Uniform quantization methods encompass Post-Training Quantization (PTQ) [30] and Quantization Aware Training (QAT). PTQ has gained popularity for its ability to quantize models without requiring retraining. However, it suffers from accuracy degradation due to its heavy reliance on input data for setting quantization parameters, leading to poor alignment between parameters and weights. QAT addresses this issue by training with inserted fake quantization nodes [16], allowing network weights to adapt to quantization errors and enhancing the accuracy of quantized models. Early works like Dorefa [45] achieve network acceleration by quantizing activation values, weights, and gradients simultaneously. Further advancements such as PACT [6], utilize learnable parameters to clip activation values to gain more reasonable quantization intervals. LSQ [9], one of the most widely used quantization methods, introduces a novel approach to estimate and scale the task loss gradient on the quantizer step size for each weight and activation layer, enabling the quantization step size to be learned with network optimization and achieving commendable results across various tasks. Given the satisfactory accuracy and notable acceleration achieved by quantized networks, quantization-based model compression has been extensively studied in various domains [35; 39; 23], demonstrating the efficacy of quantization schemes.
3 Method
3.1 Preliminaries
Gait Recognition Network Architecture
To strike a balance between universality and effectiveness, among numerous methods, we opt for GaitBase as the baseline model for compression. For appearance-based methods, given a set of captured RGB images , the input of the models are designed as binary silhouette sequences , where represents the number of frames in the sequence and is the number of data points. For GaitBase, the model can be formulated as:
where denotes the embeddings, is the separate fully connected layer, is the temporal pooling and horizontal pooling layer, is the backbone network, and denotes the connection among network parts. Given the BNNecks module , is the logit vector.
Quantization Framework
For the quantized gait recognition network, the weight and activation values of computing units (such as convolutional and linear layers) are compressed to low bit-widths by the following point-wise quantization functions:
where denotes either weights or activations of a specific layer, returns with values below set to and values above set to ( and are scopes set based on quantization type and bit-width), is a learnable parameter that adjusts the quantization step size and is the round function. For example, given a quantization bit-width , and is set to and respectively for unsigned input and and for signed input. Since the round function is non-differentiable, the derivative of with respect to during the backward pass [9; 6; 45] can be represented as:
Overall Objective
For QAT, global optimization learning algorithms optimize a suitable objective function by observing the training data as a whole to learn the parameters of deep networks:
where denotes the corresponding labels and represents the implicit combination of task loss and quantization error . The pipeline of QGait is shown in Fig. 2.
3.2 Training with Soft Quantizer
STE serves as a classical method for estimating gradients, proven to be simple yet effective in many tasks. For some learning-based quantization methods, using STE to estimate quantization errors has yielded promising results. This is because, during quantization-aware training, algorithms tend to minimize task loss rather than quantization error [9; 25]. However, in the context of gait recognition tasks, particularly in appearance-based approaches with binarized input, even minor numerical perturbations can lead to vastly different features. Therefore, it should be helpful to use a differentiable function to approximate the step function to enhance the performance of gait recognition tasks. For the round function, the maximum quantization error occurs at the midpoint between two adjacent approximate values. Additionally, we hope that the derivative of this function varies periodically with changes in input values. A function that meets the above requirements can be defined as follows:
where is the range of error, is set to the magnitude of gradient changes, and denotes the floor function. The derivative of with respect to during the backward pass becomes:
However, from an optimization perspective, while such a function can simulate errors effectively by setting different values of , it complicates the convergence of the network. Let be a convex function that is L-smooth. Generally, our iterations are given by , where is the step size and is a random direction. Then we have:
where is the smoothness coefficient. Under the framework of stochastic gradient descent (SGD), we can approximately consider .
Next, let’s analyze the two terms on the right-hand side of Eq. 8. For the first term, the gradient is always set to during the backward pass when using STE. In most machine learning libraries, the gradient of floor functions is set to . So we only need to consider part of the :
We can then compute the gradient of as:
Assuming the random variable follows a uniform distribution, we can compute the first term as (see Appendix A.1):
which is a function involving that is consistently less than as Fig. 3(b) shows. As increases, this term tends to approach , indicating that the optimization process will become increasingly challenging to continue. But we aim for as much descent as possible at each iteration, so STE is better for the first term. As for the second term, we can compute that:
which is consistently greater than as Fig shows. Combining Eq. 8, Eq. 11, and Eq. 12, we can see that while the soft quantizer simulates quantization error, STE is superior in terms of convergence as Fig. 3(c) shows. Therefore, during the initial training phase, STE is utilized to minimize task loss, while during the second fine-tuning phase, the soft quantizer is employed to balance quantization error and task loss.
3.3 Inter-class Distance-guided Distillation
In the preceding sections, we mitigated the impact of quantization error on prediction results by introducing the soft quantizer. However, on some complex datasets[33; 43], transitioning from 32-bit to 8-bit quantization still inevitably incurs some loss due to inherent information loss in the process of data discretization. Moreover, we observed that for the embeddings outputted by the 4-bit network, the relative distance between samples of different classes has changed. This inspires us to explore whether we can leverage information from high-bit networks to guide the optimization process of low-bit networks.
Given an input , the output embedding vector is . For with high bit and low bit , the logits vectors are denoted as and respectively. Traditional Knowledge Distillation (KD) approaches measure the distribution difference between two vectors. The probability vectors are calculated as:
where and are class indices, and is the temperature. Generally, the objective is realized by minimizing KL divergence:
For the well-refined low-bit model, we assume that the KL divergence loss is minimized, aligning the predicted probability density with that of the high-bit model. For any pair of class indices , it gives:
If we simplify the above equation and generalize it to all classes, we obtain: where is the function of standard deviation. In most practical applications, discrete quantization tends to reduce the variance of the data, as the quantization process essentially reduces the precision of the data, causing values that may be very close to each other to be merged into a single value or range. This merging reduces the differences between data points, thus typically leading to a decrease in variance. The experiments also confirm this, showing that the variance of logits decreases from to after quantization. Therefore, the aforementioned optimization process essentially simulates the numerical disparities brought about by quantization.
Inspired by the remarkable performance of triplet loss in gait recognition [29; 11], we visualize the distribution of embedding vectors from different samples as shown in Fig. 4. Both low-bit and high-bit networks can effectively cluster samples of the same class. Despite minor differences in inner-class distributions, the impact of quantization-induced numerical disparities on performance is negligible compared to noticeable changes in inter-class distances (dashed lines in Fig. 4). Based on the observation of heterogeneous distances, we decide to approach distillation from the perspective of distance distribution. Considering a mini-batch of data, the probability vectors guided by the self-similarity of distances between samples’ embeddings are defined as:
where denotes the labels of samples, indicates the Euclidean distance, and are the sample indices. The final objective function can be formulated as:
4 Experimental Results
To validate the generalization and effectiveness of our approach, we performed comprehensive experiments on the mainstream datasets and compared our approach with state-of-the-art gait methods and quantization methods. We selected representative methods Dorefa[45], PACT[6], and LSQ[9] as the baselines for quantization. QGait models in subsection 4.2 are implemented based on GaitBase[11]. Other quantized models are reported in subsection 4.4. The quantization bit-width is set to 4-bit or 8-bit, as in practical scenarios, devices often do not support other types of bits. The code implementation is based on PyTorch 2.0, and all experiments were conducted using RTX A6000s.
Methods | Bit-width | BitOPs (G) | Rank-1 (%) | Rank-5 | Rank-10 | mAP | mINP |
GaitSet [4] | 32 | 10.68 | 36.75 | 58.66 | 64.20 | 30.21 | 17.55 |
GaitPart [12] | 32 | 10.67 | 28.44 | 47.58 | 53.70 | 22.01 | 12.32 |
GaitGL [22] | 32 | 41.64 | 29.88 | 48.69 | 54.42 | 22.30 | 13.50 |
DyGait [37] | 32 | 652.22 | 66.30 | 80.80 | 86.10 | 56.40 | 37.30 |
GaitBase [11] | 32 | 118.30 | 64.40 | 81.50 | 85.80 | 55.28 | 36.73 |
Dorefa-4 [45] | w4/a4 | - | 60.00 | 78.30 | 83.20 | 50.32 | 32.66 |
PACT-4 [6] | w4/a4 | - | 59.80 | 78.00 | 82.80 | 49.56 | 31.53 |
LSQ-4 [9] | w4/a4 | - | 63.80 | 79.60 | 86.30 | 54.22 | 35.71 |
QGait-4 | w4/a4 | 1.85 | 64.60 | 81.20 | 86.20 | 54.29 | 35.76 |
Dorefa-8 [45] | w8/a8 | - | 64.30 | 80.80 | 86.10 | 54.34 | 35.87 |
PACT-8 [6] | w8/a8 | - | 63.30 | 80.50 | 85.20 | 53.41 | 34.64 |
LSQ-8 [9] | w8/a8 | - | 64.80 | 81.40 | 85.90 | 55.15 | 36.76 |
QGait-8 | w8/a8 | 7.39 | 66.50 | 81.30 | 86.20 | 55.82 | 37.05 |
4.1 Experimental Settings
Training Settings
We pre-process the original gait sequences for all the datasets following previous works [11; 37; 10] and the size of each frame is set to 64 × 44. For Gait3D[43], GREW[46] and OUMVLP[33], the number of Identities and the number of samples per identity is set to and for CASIA-B [40]. The optimizer is Adam with a learning rate of 1e-4 and momentum of 0.9 for all the datasets for a fair comparison. The frame number settings are consistent with OpenGait[11]. In the first training stage of QGait, the iterations are set to 60K, 180K, and 120Krespectively for Gait3D[43], GREW[46] and OUMVLP[33]. In the fine-tuning stage, we only conduct 1K or 2K iterations according to different datasets to avoid overfitting while ensuring a fair comparison relative to the full-precision model.
Evaluation Settings
We evaluated on the test set corresponding to the training dataset. In addition to comparing with quantized models, we also included comparisons with state-of-the-art full-precision models, including GaitSet[4], GaitPart[12], GaitGL[22], DyGait[36] and GaitBase[11]. Evaluation metrics used include Rank-n, mAP (mean Average Precision), mINP (mean Inverted Normalized Precision), and BitOPs (Bit Operations Per Second). Conventionally, given the weight of and input feature of , BitOPs of a quantized convolution layer can be calculated as . The BitOPs of each model on different datasets are calculated based on different frames (100 frames for Gait3D and OUMVLP, 120 frames for GREW), along with different model configurations.
Methods | Bit | BitOPs | Rank-1 | Rank-5 |
GaitSet [4] | 32 | 50.68 | 46.30 | 63.62 |
GaitPart [12] | 32 | 31.68 | 43.99 | 60.74 |
GaitGL [22] | 32 | 234.21 | 47.28 | 63.88 |
DyGait [37] | 32 | 867.14 | 71.40 | 83.21 |
GaitBase [11] | 32 | 141.96 | 60.12 | 75.47 |
Dorefa-4 [45] | w4/a4 | - | 53.31 | 69.68 |
PACT-4 [6] | w4/a4 | - | 52.40 | 68.72 |
LSQ-4 [9] | w4/a4 | - | 57.44 | 73.51 |
QGait-4 | w4/a4 | 2.22 | 58.52 | 74.06 |
Dorefa-8 [45] | w8/a8 | - | 60.00 | 74.93 |
PACT-8 [6] | w8/a8 | - | 59.03 | 73.62 |
LSQ-8 [9] | w8/a8 | - | 59.89 | 75.13 |
QGait-8 | w8/a8 | 8.87 | 60.60 | 75.89 |
Methods | Bit | BitOPs | Rank-1 | Rank-5 |
GaitSet [4] | 32 | 42.29 | 87.12 | 91.64 |
GaitPart [12] | 32 | 26.44 | 88.68 | 92.07 |
GaitGL [22] | 32 | 193.30 | 89.73 | 92.32 |
DyGait [37] | 32 | 829.79 | 90.47 | 92.89 |
GaitBase [11] | 32 | 118.30 | 90.32 | 92.67 |
Dorefa-4 [45] | w4/a4 | - | 88.24 | 92.02 |
PACT-4 [6] | w4/a4 | - | 87.65 | 91.96 |
LSQ-4 [9] | w4/a4 | - | 86.19 | 91.19 |
QGait-4 | w4/a4 | 1.85 | 89.24 | 92.29 |
Dorefa-8 [45] | w8/a8 | - | 90.26 | 92.60 |
PACT-8 [6] | w8/a8 | - | 90.13 | 92.50 |
LSQ-8 [9] | w8/a8 | - | 90.28 | 92.61 |
QGait-8 | w8/a8 | 7.39 | 90.35 | 92.63 |
4.2 Comparison with State-of-the-art Methods
Evaluation on Gait3D
The Gait3D dataset is a large-scale dataset, which contains 4,000 subjects and over 25,000 sequences captured from an unconstrained indoor scene by 39 cameras. The diversity and complexity of the data make many models perform poorly on this dataset. As Tab. 1 shows, in the 32-bit full-precision model, DyGait achieves state-of-the-art performance, albeit at the expense of costly BitOPs. While GaitSet and GaitPart have computational advantages, their accuracy falls short of requirements. QGait achieves performance over the full-precision model with very low BitOPs and achieves precision lossless compared to other quantization methods. The QGait-8 outperforms GaitBase with 2.1% increase of Rank-1 accuracy. QGait-4 surpasses other quantization methods and even outperforms the full-precision model in Rank-1 accuracy. For complex datasets like Gait3D, the introduction of quantization may act as a form of regularization, enhancing the network’s generalization ability, thus achieving significant performance improvement.
Evaluation on GREW
The GREW dataset is a large-scale outdoor gait dataset. It can be observed from Tab. 3(a) that GaitGL, despite its high computational costs (with BitOPs exceeding 200G), does not achieve the expected high performance. For 8-bit quantized models, only QGait achieves lossless compression relative to the full-precision model. While QGait surpasses other quantization methods at 4-bit, e.g., QGait achieves a Rank-1 accuracy exceeding LSQ by 1.08%, it still incurs a loss in accuracy. This may be because exploration of large-scale outdoor datasets like GREW is still insufficient in the full-precision model, thus the errors introduced by quantization cannot be effectively compensated for through training.
Evaluation on OUMVLP
Although the OUMVLP dataset contains a large amount of data, the gait sequences in this dataset are all in normal states without occlusions, disguises, or other variations. From Tab. 3(b), it can be observed that QGait incurs some loss in accuracy at 4-bit, while there is minimal loss at 8-bit. This suggests that using stronger baseline models may lead to better performance on the OUMVLP dataset. Nevertheless, compared to other quantization methods, we still maintain a lead of over 1% in accuracy at 4-bit.
4.3 Ablation Study
Strategy | T=3 | ratio | T=5 | ratio | STE |
FIXED | 62.72 | - | 62.02 | - | 64.80 |
GROW-1 | 64.88 | 0.1/100 | 65.48 | 0.1/100 | - |
GROW-2 | 65.24 | 0.2/100 | 65.89 | 0.2/100 | - |
GROW-3 | 61.15 | 1/1000 | 58.92 | 1/1000 | - |
Baseline | Vanilla KD | IDD | FT | Rank-1 (%) |
✓ | 64.80 | |||
✓ | ✓ | 64.21 | ||
✓ | ✓ | 65.64 | ||
✓ | ✓ | ✓ | 66.50 |
Training with Soft Quantizer We empirically demonstrate that introducing a soft quantizer at the beginning of training indeed significantly affects the model’s convergence. As the parameter k increases, the soft quantizer approaches the round function more closely, resulting in smaller updates to gradients at each step, as illustrated in Fig.5 in Appendix A.2. We also experimentally determine how to update the value during the fine-tuning stage. After trying several different strategies, we decided to use a gradually increasing approach, where we increment the value of k as the number of iterations increases until it reaches a predetermined threshold as Tab. 4(a). The selection of the threshold is based on the results calculated in the methodology section.
Methods | R1-NM | R1-BG | R1-CL |
GaitSet [4] | 95.79 | 89.58 | 73.62 |
GaitSet-4bit | 95.12 | 88.65 | 72.33 |
GaitSet-8bit | 95.73 | 89.44 | 73.58 |
GaitPart [12] | 96.14 | 90.69 | 78.73 |
GaitPart-4bit | 95.89 | 89.99 | 78.01 |
GaitPart-8bit | 96.15 | 90.77 | 78.92 |
DyGait [37] | 98.52 | 96.13 | 87.71 |
DyGait-4bit | 98.72 | 96.65 | 87.90 |
DyGait-8bit | 98.53 | 96.12 | 87.73 |
Inter-class Distance-guided Distillation To study the effects of IDD together with other components, we conducted module ablation experiments on the baseline method LSQ. It can be observed from Tab. 4(b) that, as discussed in the methodology section, vanilla KD overly balances the quantization error, resulting in decreased accuracy. IDD improves the accuracy by 1.84%, and even after incorporating the soft quantizer fine-tuning strategy, it still achieves a further 0.86% improvement in accuracy.
4.4 Quantization of Different Models
We also conduct experiments on different model architectures using our proposed QGait method, including GaitSet, GaitPart and DyGait. It can be seen from Tab. 4 that QGait shows comparable performance with full-precision models on different model structures. For two lightweight models, the 4-bit quantized models lost some precision. However, the quantized 4-bit DyGait model outperforms the full-precision model. That is because the quantization acts as a regularization. It can alleviate overfitting on some heavy models while the data is relatively simple. This property may inspire the later works.
5 Conclusion
In this paper, we explore for the first time a gait recognition network based on quantization compression. Addressing the data characteristics of gait recognition tasks and the difficulty of training convergence, we design quantization schemes that are different from conventional tasks. Experimental results demonstrate that the two proposed methods effectively improve the quantization accuracy of gait recognition. In experiments, we found that different datasets exhibit varying levels of robustness to quantization, which may be related to the complexity of the data. This brings new perspectives and considerations for future research in gait recognition.
Appendix A Appendix
A.1 Calculation of Integral
The methods section of this paper involves the calculation of two key integrals. The following is the derivation of the corresponding two indefinite integrals.
Let be the integral variable, the integral of can be calculated as:
For the integrand , substitute and , we get:
For the integrand , substitute and let , then we have:
where is a constant. Substitute back for and , we get:
A.2 Direct Training with Soft Quantizer
The parameter controls the steepness of the curve. As increases, the curve becomes closer to the curve of the round function. In the body part, We have demonstrated through theoretical analysis that the convergence of is worse than using STE. Here, we use different values as parameters for the soft quantizer, which are actually applied to network training. The results are as Fig. 5 shows. It is obvious that as increases, the convergence speed of the network becomes very slow. Although the soft quantizer simulates quantization errors, the decrease in gradient mean and the increase in variance make learning task losses exceptionally difficult.
A.3 Limitations
While quantization can bring actual acceleration effect to gait recognition networks, there are also some limitations:
Model Deployment: In the training stage, the model simulates the quantization error with pure floating-point numbers so as to obtain the corresponding network weight which can adapt to the error. In the inference phase, it is necessary to select the appropriate strategy for deployment according to the different deployment platforms (such as GPU and mobile). This process may involve model optimization, model transformation, model quantization, and model compilation optimization. Maintaining alignment of model accuracy during this process requires considerable experience.
Quantization of Different Models: The structural differences between different models can be significant. Quantization Aware Training on these structures requires manual modifications to the model files. At the same time, the influence of different topologies on quantization is unknown. The scalability of quantization needs to be based on the understanding of model properties and the adjustment of quantization parameters.
Hardware Support: Quantization requires specific hardware support to achieve acceleration. For example, the NVIDIA Tesla V100 graphics card supports FP32, FP16, INT8 and INT4 precision. The A100 graphics card in addition to supporting FP32 and FP16, also introduces new precision TF32 and FP64. The A100 also supports mixed precision BF16/FP16 as well as INT8, INT4, and Binary.
A.4 Broader Impacts
The quantization and compression of gait recognition networks have the potential to bring about various societal impacts, both positive and negative. Here is a detailed discussion of these impacts:
Increased Portability and Usability: Quantization and compression enable gait recognition models to be deployed on resource-constrained devices such as smartphones and IoT devices, facilitating the use of gait recognition technology in mobile and embedded systems. Reduced model size and computational requirements make real-time processing feasible, enhancing the responsiveness and user experience of gait recognition systems.
Energy Savings and Cost Reduction: Compressed models demand less computational power, thereby reducing energy consumption, which is particularly beneficial for data centers and edge devices, contributing to a greener computing environment. Lower hardware and maintenance costs due to reduced computational demands make gait recognition technology more affordable for businesses and consumers, promoting wider adoption.
Enhanced Privacy Protection: Quantized and compressed models are well-suited for running on local devices, which minimizes the need to transmit data to the cloud. This local processing helps protect sensitive information as data does not leave the device.
However, there are also potential negative impacts:
Privacy and Security Concerns: The misuse of gait recognition technology can lead to privacy infringements. Governments, companies, or individuals could potentially exploit this technology for extensive surveillance and unauthorized monitoring. As the technology becomes more ubiquitous due to easier deployment via quantized models, the risk of unauthorized data collection and analysis increases, raising privacy concerns.
Technical Limitations and Reliability Issues: Despite efforts to maintain accuracy, quantized models might suffer from reduced precision, particularly in complex scenarios or when dealing with diverse populations, which can lead to misidentification or non-recognition. Compressed models may exhibit lower adaptability to new data or varied environments, necessitating frequent updates and optimizations.
Social and Ethical Issues: The widespread application of gait recognition raises ethical questions related to biometric technology, such as transparency, informed consent, and user rights. These issues might be exacerbated with the proliferation of quantized models due to lower barriers to entry. Additionally, if the training datasets are biased, gait recognition systems could inherit these biases, leading to inconsistent recognition rates across different groups, thus causing discrimination issues.