### v2

# QGait: Toward Accurate Quantization for Gait Recognition

Abstract
Existing deep learning methods have made significant progress in gait recognition. Quantization can facilitate the application of gait models as a model-agnostic general compression technique. Typically, appearance-based models binarize inputs into silhouette sequences. However, mainstream quantization methods prioritize minimizing task loss over quantization error, which is detrimental to gait recognition with binarized inputs. To address this, we propose a differentiable soft quantizer, which better simulates the gradient of the round function during backpropagation. This enables the network to learn from subtle input perturbations. However, our theoretical analysis and empirical studies reveal that directly applying the soft quantizer can hinder network convergence. We addressed this issue by adopting a two-stage training strategy, introducing a soft quantizer during the fine-tuning phase. However, in the first stage of training, we observed a significant change in the output distribution of different samples in the feature space compared to the full-precision network. It is this change that led to a loss in performance. Based on this, we propose an Inter-class Distance-guided Calibration (IDC) strategy to preserve the relative distance between the embeddings of samples with different labels. Extensive experiments validate the effectiveness of our approach, demonstrating state-of-the-art accuracy across various settings and datasets. Code is available at link.
1 Introduction
In recent years, gait recognition has emerged as a prominent field due to its potential in long-range pedestrian retrieval based on walking patterns. Unlike other biometric technologies, gait, as a form of locomotion, is difficult to disguise [Lam2007HumanGR, Connor2018BiometricRB]. Additionally, gait recognition is robust against common covariates such as attire, carrying items, and standing conditions [Guan2015OnRT, Choudhury2016ClothingAC]. Therefore, gait recognition methods hold significant promise for capturing biometric features remotely in uncontrolled environments, thus offering broad application prospects.
With the development of deep learning, gait recognition methods based on deep features [Wu2017ACS, Makihara2017JointIA, Fan2020GaitPartTP, Hou2020GaitLN, Lin2020GaitRV, Lin2020GaitRW, guo2025gait, wang2025gaitadapt] have achieved notable results on various datasets [Yu2006AFF, Takemura2018MultiviewLP, Zheng2022GaitRI, Shen2022LidarGaitB3, Zhu2021GaitRI, Fan2022OpenGaitRG, Fan2023ExploringDM]. Currently, state-of-the-art gait recognition technologies can achieve recognition distances exceeding hundreds of meters even with 4K high-definition cameras [Shen2022LidarGaitB3]. However, existing methods have scarcely considered the increasing computational resources required by the advanced edge devices. In real-world gait recognition scenarios, especially when real-time processing of video streams is required, the inference overhead of gait models cannot be overlooked as Fig. 1 shows. While methods based on CNNs [Fan2022OpenGaitRG, Zheng2022GaitRI] or Transformers [Fan2023ExploringDM] have demonstrated excellent performance, edge devices in deployment environments cannot meet the demands of these resource-intensive algorithms. This necessitates reducing the memory and computational burden of gait recognition methods while preserving performance, enabling deployment on resource-constrained devices. Current efforts to compress and accelerate neural networks primarily focus on quantization [Zhou2016DoReFaNetTL, Choi2018PACTPC, Esser2019LearnedSS, Lee2021NetworkQW] and pruning [Lin2020HRankFP, Zhou2021LearningNM, He2018FilterPV]. Compared to pruning, quantization is more deployment-friendly and offers significant acceleration, particularly Quantization Aware Training [Jacob2017QuantizationAT].
Existing quantization methods [Zhang2018LQNetsLQ, Baskin2018NICENI, Esser2019LearnedSS] have achieved decent results on some tasks, but applying them to gait recognition tasks has failed to deliver comparable performance. These methods use a Straight-Through Estimator (STE) [Bengio2013EstimatingOP] to approximate the gradient of non-differentiable operations during the backward pass. While STE can smoothly transfer gradients, it is not sensitive to decimal-level changes. This may have no impact on other tasks but can be fatal for changes in gait posture. For example, in RGB input, the variation from 0 to 1 in a small area is not significant. However, for binarized gait silhouettes, this may indicate a change in posture. In this paper, we propose using parameterized soft quantization operators to prevent the network from accumulating quantization errors. Initially, we attempted to directly utilize a soft quantizer in the training process as a replacement for STE. However, through theoretical analysis and empirical validation, we have demonstrated that although this approach can simulate the quantization error introduced by non-differentiable operations, it significantly impacts the network’s convergence. To harness the strengths of both STE and the soft quantizer while mitigating their respective limitations, we propose a two-stage training strategy. In the first stage, we employ STE to facilitate the network’s convergence towards the vicinity of the optimal solution. Subsequently, we utilize the soft quantizer in the fine-tuning stage, enabling the simulation of decimal-level errors, which improves the network performance.
After quantization, the 8-bit models exhibit outstanding performance, surpassing even the full-precision model. However, when quantized to lower bit widths, the model’s accuracy experiences significant degradation. Through analysis of the sample distribution in feature space, we observed significant changes in the relative positions of samples between the 4-bit quantized model and the full-precision model. This change could be a significant contributing factor to the observed decrease in accuracy. This insight prompts us to consider whether a calibration strategy could be employed to align the sample distribution of the low-bit models to high-bit models. A possible alignment solution is to distill the distribution from high-bit models. However, through an examination of the properties of naive distillation [Esser2019LearnedSS, Kim2019QKDQK, Yao2022ZeroQuantEA], we found that such methods result in the low-bit model learning the numerical discrepancies between the outputs of the full-precision model and itself, which are discrepancies inherently introduced by quantization. Therefore, we propose an Inter-class Distance-guided Calibration strategy to focus more on the differences in sample distribution between classes rather than the numerical differences within classes.
Our contributions can be summarized as follows:
We introduce a quantization-based gait recognition network (QGait), which performs lossless compression and acceleration on existing methods as Fig. 2 shows.
We propose a two-stage quantization training strategy together with the IDC method, which simultaneously optimizes quantization error and task loss.
We conduct detailed experiments and the results demonstrate that our proposed QGait achieves comparable performance to full-precision networks with the lowest computational overhead.
2 Related Work
2.1 Gait Recognition
Gait recognition methods can generally be classified into two categories: model-based methods [Li2020EndtoEndMG, Liao2017PoseBasedTN] and appearance-based methods [Wang2023DyGaitED, Shen2022GaitRW]. Model-based approaches tend to utilize estimated human body models as input, such as 2D/3D poses and SMPL [Loper2023SMPLAS]. For instance, GaitGraph [Teepe2021GaitgraphGC] employs graph convolutional networks for learning gait representations based on 2D skeletons, SMPLGait [Zheng2022GaitRI] leverages the 3D geometric information of SMPL models to enhance the learning of gait appearance features, and GaitTR [Zhang2022SpatialTN] combines transformer and convolutional layers to represent spatial and temporal information respectively. However, although model-based methods theoretically exhibit robustness to factors like carrying items and clothing, they often perform poorly due to the error accumulation in the pre-processing stage while entailing high computational costs, thereby potentially lacking practicality in certain real-world scenarios. On the other hand, appearance-based methods directly learn shape features from input videos, offering simplicity and ease of use while preserving privacy. With the rise of deep learning, most current appearance-based works focus on spatial feature extraction and gait temporal modeling. Specifically, GaitSet [Chao2018GaitSetRG] treats gait sequences as a set for the first time, employing the maximum function to compress sequences of frame-level spatial features. GaitPart [Fan2020GaitPartTP] meticulously explores minor differences in various parts of the silhouette to model periodic gait features. GaitGL [Lin2020GaitRV] addresses the limitations of spatially global gait representations in neglecting details, and locally based descriptors in failing to capture relationships between adjacent parts, by developing global and local convolutional layers. Building upon these works, GaitBase [Fan2022OpenGaitRG] significantly simplifies gait modeling and achieves excellent results with its simple yet effective design. Considering that existing methods often lack a practical perspective, it becomes necessary to design a lightweight network to aid gait recognition models in real-world applications. To the best of our knowledge, the proposed QGait is the first quantization-based compressed gait recognition network.
2.2 Network Quantization
Quantization [Chen2024AdaptiveQW, Cai2023BinarizedSC, Rastegari2016XNORNetIC] is a widely adopted technique in computer vision for compressing and accelerating neural networks. The advantage of quantization is that it can be easily combined with pruning, distillation, and other methods to further compress the network. Depending on the mapping strategy, quantization can be categorized into uniform and non-uniform [Liu2021NonuniformtoUniformQT] quantization. Non-uniform quantization necessitates specific hardware support, thus current research predominantly focuses on uniform quantization exploration. Uniform quantization methods encompass Post-Training Quantization (PTQ) [Shang2022PostTrainingQO] and Quantization Aware Training (QAT). PTQ has gained popularity for its ability to quantize models without requiring retraining. However, it suffers from accuracy degradation due to its heavy reliance on input data for setting quantization parameters, leading to poor alignment between parameters and weights. QAT addresses this issue by training with inserted fake quantization nodes [Jacob2017QuantizationAT], allowing network weights to adapt to quantization errors and enhancing the accuracy of quantized models. Early works like Dorefa [Zhou2016DoReFaNetTL] achieve network acceleration by quantizing activation values, weights, and gradients simultaneously. Further advancements such as PACT [Choi2018PACTPC], utilize learnable parameters to clip activation values to gain more reasonable quantization intervals. LSQ [Esser2019LearnedSS], one of the most widely used quantization methods, introduces a novel approach to estimate and scale the task loss gradient on the quantizer step size for each weight and activation layer, enabling the quantization step size to be learned with network optimization and achieving commendable results across various tasks. Given the satisfactory accuracy and notable acceleration achieved by quantized networks, quantization-based model compression has been extensively studied in various domains [Tian2023CABMCB, Yao2022ZeroQuantEA, Lin2023AWQAW], demonstrating the efficacy of quantization schemes. However, the impact of quantization in gait recognition remains underexplored. In this paper, we introduce QGait, which, built upon optimized quantization training strategies and tailored to the characteristics of gait recognition tasks, unveils the effectiveness of quantization in compressing gait recognition networks for the first time, offering novel perspectives and insights into gait recognition research.
3 Method
3.1 Preliminaries
Gait Recognition Network Architecture
To strike a balance between universality and effectiveness, among numerous methods, we opt for GaitBase as the baseline model for compression. For appearance-based methods, given a set of captured RGB images , the input of the models are designed as binary silhouette sequences , where represents the number of frames in the sequence and is the number of data points. For GaitBase, the model can be formulated as:
where denotes the embeddings, is the separate fully connected layer, is the temporal pooling and horizontal pooling layer, is the backbone network, and denotes the connection among network parts. Given the BNNecks module , is the logits vector.
Quantization Framework
For the quantized gait recognition network, the weight and activation values of computing units (such as convolutional and linear layers) are compressed to low bit-widths by the following point-wise quantization functions:
where denotes either weights or activations of a specific layer, returns with values below set to and values above set to ( and are scopes set based on quantization type and bit-width), is a learnable parameter that adjusts the quantization step size and is the round function. For example, given a quantization bit-width , and is set to and respectively for unsigned input and and for signed input. Since the round function is non-differentiable, the derivative of with respect to during the backward pass [Esser2019LearnedSS, Choi2018PACTPC, Zhou2016DoReFaNetTL] can be represented as:
Overall Objective
For QAT, global optimization learning algorithms optimize a suitable objective function by observing the training data as a whole to learn the parameters of deep networks:
where denotes the corresponding labels and represents the implicit combination of task loss and quantization error . The pipeline of QGait is shown in Fig. 3.
3.2 Training with Soft Quantizer
STE serves as a classical method for estimating gradients, proven to be simple yet effective in many tasks. For some learning-based quantization methods, using STE to estimate quantization errors has yielded promising results. This is because, during quantization-aware training, algorithms tend to minimize task loss rather than quantization error [Esser2019LearnedSS, Liu2021NonuniformtoUniformQT]. However, in the context of gait recognition tasks, particularly in appearance-based approaches with binarized input, even minor numerical perturbations can lead to vastly different features. Therefore, it should be helpful to use a differentiable function to approximate the step function to enhance the performance of gait recognition tasks. For the round function, the maximum quantization error occurs at the midpoint between two adjacent approximate values. Additionally, we hope that the derivative of this function varies periodically with changes in input values. A function that meets the above requirements can be defined as follows:
where is the range of error, is set to the magnitude of gradient changes, and denotes the floor function. The derivative of with respect to during the backward pass becomes:
However, from an optimization perspective, while such a function can simulate errors effectively by setting different values of , it complicates the convergence of the network. Let be a convex function that is L-smooth. Generally, our iterations are given by , where is the step size and is a random direction. Then we have:
where is the smoothness coefficient. Under the framework of stochastic gradient descent (SGD), we can approximately consider .
Next, let’s analyze the two terms on the right-hand side of Eq. 8. For the first term, the gradient is always set to during the backward pass when using STE. In most machine learning libraries, the gradient of floor functions is set to . So we only need to consider part of the :
We can then compute the gradient of as:
Assuming the random variable follows a uniform distribution, we can compute the first term as:
which is a function involving that is consistently less than as Fig. 4(b) shows. As increases, this term tends to approach , indicating that the optimization process will become increasingly challenging to continue. However, we aim for as much descent as possible at each iteration, so STE is better for the first term. As for the second term, we can compute that:
which is consistently greater than . Combining Eq. 8, Eq. 11, and Eq. 12, we can see that while the soft quantizer simulates quantization error, STE is superior in terms of convergence as Fig. 4(c) shows. Therefore, during the initial training phase, STE is utilized to minimize task loss, while during the second fine-tuning phase, the soft quantizer is employed to balance quantization error and task loss.
3.3 Inter-class Distance-guided Calibration
| Methods | Bit-width | BitOPs (G) | Rank-1 (%) | Rank-5 | Rank-10 | mAP | mINP |
| GaitSet [Chao2018GaitSetRG] | 32 | 10.68 | 36.75 | 58.66 | 64.20 | 30.21 | 17.55 |
| GaitPart [Fan2020GaitPartTP] | 32 | 10.67 | 28.44 | 47.58 | 53.70 | 22.01 | 12.32 |
| GaitGL [Lin2020GaitRV] | 32 | 41.64 | 29.88 | 48.69 | 54.42 | 22.30 | 13.50 |
| DyGait [Wang2023DyGaitED] | 32 | 652.22 | 66.30 | 80.80 | 86.10 | 56.40 | 37.30 |
| GaitBase [Fan2022OpenGaitRG] | 32 | 118.30 | 64.40 | 81.50 | 85.80 | 55.28 | 36.73 |
| Dorefa-4 [Zhou2016DoReFaNetTL] | w4/a4 | - | 60.00 | 78.30 | 83.20 | 50.32 | 32.66 |
| PACT-4 [Choi2018PACTPC] | w4/a4 | - | 59.80 | 78.00 | 82.80 | 49.56 | 31.53 |
| LSQ-4 [Esser2019LearnedSS] | w4/a4 | - | 63.80 | 79.60 | 86.30 | 54.22 | 35.71 |
| QGait-4 | w4/a4 | 1.85 | 64.60 | 81.20 | 86.20 | 54.29 | 35.76 |
| Dorefa-8 [Zhou2016DoReFaNetTL] | w8/a8 | - | 64.30 | 80.80 | 86.10 | 54.34 | 35.87 |
| PACT-8 [Choi2018PACTPC] | w8/a8 | - | 63.30 | 80.50 | 85.20 | 53.41 | 34.64 |
| LSQ-8 [Esser2019LearnedSS] | w8/a8 | - | 64.80 | 81.40 | 85.90 | 55.15 | 36.76 |
| QGait-8 | w8/a8 | 7.39 | 66.50 | 81.30 | 86.20 | 55.82 | 37.05 |
In the preceding sections, we mitigated the impact of quantization error on prediction results by introducing the soft quantizer. However, on some complex datasets [Takemura2018MultiviewLP, Zheng2022GaitRI], transitioning from 32-bit to 8-bit quantization still inevitably incurs some loss due to inherent information loss in the process of data discretization. Moreover, we observed that for the embeddings outputted by the 4-bit network, the relative distance between samples of different classes has changed. This inspires us to explore whether we can leverage information from high-bit networks to guide the optimization process of low-bit networks.
Given an input , the output embedding vector is . For with high bit and low bit , the logits vectors are denoted as and respectively. Naive Knowledge Distillation (KD) approaches measure the distribution difference between two vectors. The probability vectors are calculated as:
where and are class indices, and is the temperature. Generally, the objective is realized by minimizing KL divergence:
For the well-trained low-bit model, we assume that the KL divergence loss is minimized, aligning the predicted probability density with that of the high-bit model. For any pair of class indices , it gives:
If we simplify the above equation and generalize it to all classes, we obtain: where is the function of standard deviation. In most practical applications, discrete quantization tends to reduce the variance of the data, as the quantization process essentially reduces the precision of the data, causing values that may be very close to each other to be merged into a single value or range. This merging reduces the differences between data points, thus typically leading to a decrease in variance. The experiments also confirm this, showing that the variance of logits decreases from to after quantization. Therefore, the aforementioned optimization process essentially simulates the numerical disparities brought about by quantization.
Inspired by the remarkable performance of triplet loss in gait recognition [Schroff2015FaceNetAU, Fan2022OpenGaitRG], we visualize the distribution of embedding vectors from different samples as shown in Fig. 5. Both low-bit and high-bit networks can effectively cluster samples of the same class. Despite minor differences in inner-class distributions, the impact of quantization-induced numerical disparities on performance is negligible compared to noticeable changes in inter-class distances (dashed lines in Fig. 5). Based on the observation of heterogeneous distances, we decided to approach a calibration from the perspective of distance distribution. Considering a mini-batch of data, the probability vectors guided by the self-similarity of distances between samples’ embeddings are defined as:
where denotes the labels of samples, indicates the Euclidean distance, and are the sample indices. The final objective function can be formulated as:
4 Experimental Results
To validate the generalization and effectiveness of our approach, we performed comprehensive experiments on the mainstream datasets and compared our approach with state-of-the-art gait methods and quantization methods. We selected representative methods Dorefa [Zhou2016DoReFaNetTL], PACT [Choi2018PACTPC], and LSQ [Esser2019LearnedSS] as the baselines for quantization. QGait models in subsection 4.2 are implemented based on GaitBase [Fan2022OpenGaitRG]. Other quantized models are reported in subsection 4.4. The code implementation is based on PyTorch 2.0, and all experiments were conducted using RTX A6000s.
| Methods | Bit | BitOPs | Rank-1 | Rank-5 |
| GaitSet | 32 | 50.68 | 46.30 | 63.62 |
| GaitPart | 32 | 31.68 | 43.99 | 60.74 |
| GaitGL | 32 | 234.21 | 47.28 | 63.88 |
| DyGait | 32 | 867.14 | 71.40 | 83.21 |
| GaitBase | 32 | 141.96 | 60.12 | 75.47 |
| Dorefa-4 | w4/a4 | - | 53.31 | 69.68 |
| PACT-4 | w4/a4 | - | 52.40 | 68.72 |
| LSQ-4 | w4/a4 | - | 57.44 | 73.51 |
| QGait-4 | w4/a4 | 2.22 | 58.52 | 74.06 |
| Dorefa-8 | w8/a8 | - | 60.00 | 74.93 |
| PACT-8 | w8/a8 | - | 59.03 | 73.62 |
| LSQ-8 | w8/a8 | - | 59.89 | 75.13 |
| QGait-8 | w8/a8 | 8.87 | 60.60 | 75.89 |
4.1 Experimental Settings
Training Settings
We pre-process the original gait sequences for all the datasets following previous works [Fan2022OpenGaitRG, Wang2023DyGaitED, Fan2023ExploringDM] and the size of each frame is set to 64 × 44. For Gait3D [Zheng2022GaitRI], GREW [Zhu2021GaitRI] and OUMVLP [Takemura2018MultiviewLP], the number of Identities and the number of samples per identity is set to and for CASIA-B [Yu2006AFF]. The optimizer is Adam with a learning rate of 1e-4 and momentum of 0.9 for all the datasets for a fair comparison. The frame number settings are consistent with OpenGait [Fan2022OpenGaitRG]. In the first training stage of QGait, the iterations are set to 60K, 180K, and 120Krespectively for Gait3D [Zheng2022GaitRI], GREW [Zhu2021GaitRI] and OUMVLP [Takemura2018MultiviewLP]. In the fine-tuning stage, we only conduct 1K or 2K iterations according to different datasets to avoid overfitting while ensuring a fair comparison relative to the full-precision model.
Evaluation Settings
We conducted the evaluation on the test set corresponding to the training dataset. In addition to comparing with quantized models, we also included comparisons with state-of-the-art full-precision models, including GaitSet [Chao2018GaitSetRG], GaitPart [Fan2020GaitPartTP], GaitGL [Lin2020GaitRV], DyGait [Wagg2004OnAM] and GaitBase [Fan2022OpenGaitRG]. Evaluation metrics used include Rank-n, mAP (mean Average Precision), mINP (mean Inverted Normalized Precision), and BitOPs (Bit Operations Per Second). Conventionally, given the weight of -bit and input feature of -bit, BitOPs of a quantized convolution layer can be calculated as . The BitOPs of each model on different datasets are calculated based on different frames (100 frames for Gait3D and OUMVLP, 120 frames for GREW), along with different model configurations.
| Methods | Bit | BitOPs | Rank-1 | Rank-5 |
| GaitSet | 32 | 42.29 | 87.12 | 91.64 |
| GaitPart | 32 | 26.44 | 88.68 | 92.07 |
| GaitGL | 32 | 193.30 | 89.73 | 92.32 |
| DyGait | 32 | 829.79 | 90.47 | 92.89 |
| GaitBase | 32 | 118.30 | 90.32 | 92.67 |
| Dorefa-4 | w4/a4 | - | 88.24 | 92.02 |
| PACT-4 | w4/a4 | - | 87.65 | 91.96 |
| LSQ-4 | w4/a4 | - | 86.19 | 91.19 |
| QGait-4 | w4/a4 | 1.85 | 89.24 | 92.29 |
| Dorefa-8 | w8/a8 | - | 90.26 | 92.60 |
| PACT-8 | w8/a8 | - | 90.13 | 92.50 |
| LSQ-8 | w8/a8 | - | 90.28 | 92.61 |
| QGait-8 | w8/a8 | 7.39 | 90.35 | 92.63 |
4.2 Comparison with State-of-the-art Methods
Figure 1, Fig. 2, and Fig. 3 respectively provide comparisons between QGait and various SOTA models on different datasets. QGait achieves the optimal balance between model performance and computational complexity. For complex datasets like Gait3D, the introduction of quantization may act as a form of regularization, enhancing the network’s generalization ability.
4.3 Ablation Study
| Strategy | T=3 | ratio | T=5 | ratio | STE |
| FIXED | 62.72 | - | 62.02 | - | 64.80 |
| GROW-1 | 64.88 | 0.1/100 | 65.48 | 0.1/100 | - |
| GROW-2 | 65.24 | 0.2/100 | 65.89 | 0.2/100 | - |
| GROW-3 | 61.15 | 1/1000 | 58.92 | 1/1000 | - |
| Baseline | Vanilla KD | IDC | FT | Rank-1 (%) |
| ✓ | 64.80 | |||
| ✓ | ✓ | 64.21 | ||
| ✓ | ✓ | 65.64 | ||
| ✓ | ✓ | ✓ | 66.50 |
| Methods | R1-NM | R1-BG | R1-CL |
| GaitSet [Chao2018GaitSetRG] | 95.79 | 89.58 | 73.62 |
| GaitSet-4bit | 95.12 | 88.65 | 72.33 |
| GaitSet-8bit | 95.73 | 89.44 | 73.58 |
| GaitPart [Fan2020GaitPartTP] | 96.14 | 90.69 | 78.73 |
| GaitPart-4bit | 95.89 | 89.99 | 78.01 |
| GaitPart-8bit | 96.15 | 90.77 | 78.92 |
| DyGait [Wang2023DyGaitED] | 98.52 | 96.13 | 87.71 |
| DyGait-4bit | 98.72 | 96.65 | 87.90 |
| DyGait-8bit | 98.53 | 96.12 | 87.73 |
Training with Soft Quantizer
We empirically demonstrate that introducing a soft quantizer at the beginning of training indeed significantly affects the model’s convergence. As the parameter increases, the soft quantizer approaches the round function more closely, resulting in smaller updates to gradients at each step as Fig. 6 shows. We also experimentally determine how to update the value during the fine-tuning stage. After trying several different strategies, we decided to use a gradually increasing approach, where we increment the value of k as the number of iterations increases until it reaches a predetermined threshold as Fig. 4 shows.
Inter-class Distance-guided Calibration
To study the effects of IDC together with other components, we conducted module ablation experiments on the baseline method LSQ. It can be observed from Fig. 5 that, as discussed in the methodology section, vanilla KD overly balances the quantization error, resulting in decreased accuracy. IDC improves the accuracy by 1.84%, and even after incorporating the soft quantizer fine-tuning strategy, it still achieves a further 0.86% improvement in accuracy.
4.4 Quantization of Different Models
We also conduct experiments on different model architectures using our proposed QGait method, including GaitSet, GaitPart and DyGait. It can be seen from Fig. 6 that QGait shows comparable performance with full-precision models on different model structures. For two lightweight models, the 4-bit quantized models lost some precision. However, the quantized 4-bit DyGait model outperforms the full-precision model. That is because the quantization acts as a regularization. It can alleviate overfitting on some heavy models while the data is relatively simple. This property may inspire the later works.
4.5 Quantization of Different Backbones
Future gait models may require larger-scale backbones, rather than being limited to just ResNet9. To demonstrate the effectiveness of quantization, we individually quantized different backbones to observe their latency as Fig. 7 shows.
| Bit | Params | BitOPs | Latency | |
| ResNet-18 | 32 | 44.6M | 1858G | 18.2ms |
| ResNet-18 | 8 | 11.1M | 116G | 5.7ms |
| ResNet-18 | 4 | 5.8M | 34G | 4.1ms |
| ResNet-50 | 32 | 97.8M | 6951G | 50.6ms |
| ResNet-50 | 8 | 24.5M | 247G | 15.9ms |
| ResNet-50 | 4 | 13.1M | 67G | 11.4ms |
5 Conclusion
In this paper, we explore a gait recognition network based on quantization compression for the first time. Addressing the data characteristics of gait recognition tasks and the difficulty of training convergence, we design quantization schemes that are different from conventional tasks. Experimental results demonstrate that the two proposed methods effectively improve the quantization accuracy of gait recognition. In experiments, we found that different datasets exhibit varying levels of robustness to quantization, which may be related to the complexity of the data. This brings new perspectives and considerations for future research in gait recognition.